OpenStudy (anonymous):

Find the circumference of the circle. Use 3.14 for π and round to at least 1 decimal place.

OpenStudy (anonymous):



OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to apply this formula: circumference = 2*3.14*6=...?

OpenStudy (anonymous):

37.68

OpenStudy (er.mohd.amir):

circumference = 2 * 3.14 *6 =37.68

OpenStudy (anonymous):

okay, it sayd move one decimal pouint, so what would our answrr be than?